# Perturbation-based Effect Measures for Compositional Data

### v4

@@ Abstract
Keywords: compositional data, effect estimation, causality, perturbation, feature importance, diversity effects, nonparametrics, mixture experiments
@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982). We summarize the historical development of these fields in Section S2.1 of the supplementary material. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex. In Section 4.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Section 4.1). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function of $x$. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$. Then, for any $x\in\mathbb{R}^{d}$, we have
@@
Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined. The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$. We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures. In Section S2.2 of the supplementary material we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
A perturbation, formally defined in Section 2, is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach. By marginalizing over $Z$ the perturbation-based approach leads to the average effect $\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$ In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
@@
which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity. A key strength of using perturbations to define target parameters is that it makes them easy to communicate, interpret and discuss with practitioners. From our experience, practitioners often have a good understanding of what perturbation they are interested in and formalizing them mathematically as above is straight-forward. Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities. To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3 of the supplementary material.
@@
The paper is structured as follows. Our perturbation-based framework is formally introduced and developed in Section 2. We describe how to construct perturbations in Section 2.1 and provide a table of pre-defined simplex effects along with examples of their use in Section 2.2. In Section 3 we give an overview of the semiparametric techniques used to estimate our proposed effects. In Section 4 we provide a semi-synthetic and two real world applications of our proposal. The supplementary material contains technical details on theory and algorithms along with additional examples and experiments.
@@ 1.1 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator. All references to sections and results beginning with ‘S’ point to the supplementary material.
@@ 2 Target parameters via perturbations
We construct interpretable perturbations in Section 2.1.2. Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex. The concept of a perturbation can be defined in greater generality than we do here but we stick to the simplex for simplicity.
@@
It would be convenient to also phrase the average directional perturbation effect $\tau_{\psi}$ as a quantity that is similarly amenable to semiparametric analysis and permits us to impose a partially linear model. To motivate the upcoming condition, we consider what happens when $Z$ takes values in $\mathbb{R}^{d}$. In that case, we can write
@@
From this result it follows that if we are able to find a derivative-isolating reparametrization the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993). It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (5)
Details on the different estimation approaches are discussed in Section 3.
@@ 2.1.1 Endpoints in the simplex
2.1.2 Directions, speeds and derivative-isolating reparametrizations
@@
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed. In Section S2.7 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction in Section S2.5 to start from either a given interpretable speed or a summary statistic. We illustrate both options in the following two examples.
@@
which is scale-invariant and therefore meaningful to consider on the simplex. On the simplex, this corresponds to the speed
$s_{\psi}(z)=2z^{j}(1-z^{j})$
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
@@
Figure S1 contains an illustration of the different perturbations we have considered in this work.
2.2 Perturbation effects on the simplex
To make our proposal more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1. We motivate each perturbation effect with an example application to showcase potential use cases. In Section S2.6 we describe an additional class of perturbation effects based on general amalgamating perturbations.
2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.2, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then want to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. We already considered these perturbations in Example 1 where we examined two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed. $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S2.8). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 2 such that $L$ is equal to one minus the Gini coefficient. The construction in Example 2 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.1.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression. In Section 4.1.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
We do not define a binary centering perturbation here as it requires a positive probability of observing $Z=z_{\mathrm{cen}}$, which is rarely the case.

### v5

@@ Abstract
Keywords: feature importance, causality, nonparametrics, semiparametrics
@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982). We summarize the historical development of these fields in Section S2.1 (Sections beginning with ‘S’ refer to the supplement throughout the paper). A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex. In Section 5.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects. In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function of $x$. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$. Then, for any $x\in\mathbb{R}^{d}$, we have
@@
In the case of compositional $Z$, we will need to use perturbations that are tailored to the simplex structure of $Z$ instead of $\psi$ given above. In Section S2.2 we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
A perturbation, formally defined in Section 2, is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$, $\psi(z,1)^{j}=0$ and $\psi(z,1)^{-j}=\mathsf{C}(z^{-j})$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach. By marginalizing over $Z$ the perturbation-based approach leads to the average effect $\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$ In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
@@
which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity. A key strength of using perturbations to define target parameters is that it makes them easy to communicate, interpret and discuss with practitioners. From our experience, practitioners often have a good understanding of what perturbation they are interested in and formalizing them mathematically as above is straight-forward. Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities. To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3.
@@
The paper is structured as follows. Our perturbation-based framework is formally introduced and developed in Section 2. We describe how to construct perturbations and provide a table of pre-defined simplex effects in Section 3. In Section 4 we describe the semiparametric techniques used to estimate our proposed effects. In Section 5 we provide a semi-synthetic and two real world applications of our proposal. The supplementary material contains technical details on theory and algorithms along with additional examples and experiments.
@@ 1.1 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
@@ 2 Target parameters via perturbations
We construct interpretable perturbations in Section 3.2. Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex. The concept of a perturbation can be defined in greater generality than we do here but we stick to the simplex for simplicity.
@@
It would be convenient to also phrase the average directional perturbation effect $\tau_{\psi}$ as a quantity that is similarly amenable to semiparametric analysis and permits us to impose a partially linear model. To motivate the upcoming condition, we consider what happens when $Z$ takes values in $\mathbb{R}^{d}$. In that case, we can write (with $\omega_{\psi}$ defined in Definition 1)
@@
From this result it follows that if we are able to find a derivative-isolating reparametrization the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993). It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that $\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ Details on the different estimation approaches are discussed in Section 4.
@@ 3.1 Endpoints in the simplex
3.2 Directions, speeds and derivative-isolating reparametrizations
@@
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. To avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed. In Section S2.7 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (5) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect.
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction in Section S2.5 to start from either a given interpretable speed or a summary statistic.
3.3 Effects of changes in individual components and diversity
In this section, we give examples of how to construct directional perturbations which summarize perturbation effects of practical interest.
@@ 3.3 Effects of changes in individual components and diversity
which is scale-invariant and therefore meaningful to consider on the simplex. On the simplex, this corresponds to the speed $s_{\psi}(z)=2z^{j}(1-z^{j})$ which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
@@
Perturbation effects that capture changes in individual components, are, for example, useful as a variable selection tool in compositional data. In Section 5.2, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then want to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. We denote by $\mathrm{CFI}^{j}_{\mathrm{unit}}$ the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ the perturbation that moves at multiplicative speed. $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@
The effect of increasing diversity is often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 3.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 2 such that $L$ is equal to one minus the Gini coefficient. In Section 5.1 we give a semi-synthetic and real data example of the application of $\mathrm{CDI}_{\mathrm{Gini}}$.
In Section S2.6 we discuss average perturbation effects that capture the effect of applying amalgamations to $Z$ on $Y$ (see Section 3.1 for a definition of a general amalgamation). We summarize our predefined perturbation effects in Table 1.